# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v1

@@ 3.7. Deployment Challenges and Solution Framework (RQ4 and RQ5)
S10 Multi-cloud PaaS Deployment Overcoming challenges and high entry barriers for deploying a multi-cloud PaaS infrastructure Lightweight Proxies for PaaS Integration Usage of lightweight proxies for individual PaaS services integration
S85 IoT Application Deployment Addressing challenges related to deploying IoT applications in multi-cloud environment DRA Framework and CI Broker for Multi-Cloud Deployment Utilization of the DRA framework and CI broker component for multi-cloud deployment
@@
S17 Legacy Code Migration Deployment Handling legacy code migration, tenant engine separation, and composition modeling during container deployment in multi-cloud environment Reusable Architectural Pattern with Docker and Cloudant Adoption of a reusable architectural pattern with Docker containers and Cloudant persistence layer
S23 Tenant Performance Deployment Managing tenant service performance competition during multi-cloud container deployment Multi-Tenant and Multi-Instance Hybrid Deployment Implementation of a multi-tenant and multi-instance hybrid deployment scheme based on container technology
S45 Monitoring Deployment Overcoming the limitations of current monitoring tools during multi-cloud container deployment Platform-Independent Monitoring System Implementation of JCatascopia, a platform-independent and interoperable monitoring system
@@
S28 Autonomic System Deployment Implementing an autonomic management system for deploying multi-layer, multi-tier web architectures in multi-cloud environment Self-Tuning Performance Model and Autonomic Management Development of a self-tuning performance model and autonomic management system
S46 Microservice Deployment Managing the complexity of deploying microservice architectures in multi-cloud environment Metrics and Requirements Framework for Microservices Framework for establishing requirements and relevant metrics for the deployment of microservice-based applications
S76 Multi-cloud Application Configuration Deployment Overcoming challenges related to configuring applications during multi-cloud deployment Runtime Environment with Object Store and Artifact Repository Deployment of a runtime environment with an object store service and artifact repository
S80 Elastic Container Platform Dependency Deployment Addressing vendor lock-in and dependency on specific elastic container platforms during multi-cloud deployment Separation of Elastic Platform and Cloud Application Separation of elastic platform definition from cloud application definition
@@
S35 Inter-edge Bandwidth Deployment Addressing challenges related to interconnecting distributed localities, provisioning inter-edge bandwidth, and centralized edge nodes during multi-cloud deployment Modular Edge Cloud Computing Architecture Designing a modular edge cloud computing system architecture using containerization techniques
S41 Network Latency Deployment Minimizing the impact of network latency on application performance during multi-cloud deployment VM Selection for Composite Applications Selection of proper VMs from the global cloud marketplace to support multiple composite applications
S43 Latency Reduction Deployment Tackling the issue of high end-to-end latencies during multi-cloud deployment Edge-Cellular Hybrid Infrastructure Provisioning Provisioning infrastructure at the edge of the cellular network combined with traditional cloud infrastructures
S63 Vendor Lock-in Prevention Deployment Preventing negative effects of vendor lock-in during multi-cloud container deployment Interoperability Layer Above Cloud Infrastructure Implementation of a layer above the cloud service provider’s infrastructure for interoperability
@@
S08 ACS Compatibility Deployment Ensuring compatibility with existing ACSs and cloud solutions during container deployment in multi-cloud environment Scalable Multi-Cloud Deployment Framework Proposal and development of a scalable deployment framework for multi-cloud applications
S30 Manual Intervention Risk Deployment Managing risks related to manual intervention, vulnerability, and exposed services during multi-cloud deployment Automated Security Measures at Deployment Automated implementation of required security measures at the point of deployment
S44 Multi-provider Deployment Handling the challenges of using multiple providers and technologies during multi-cloud container deployment TOSCA-Based Deployment Modeling Approach Adoption of a TOSCA-based deployment modeling approach
S81 Cloud Provider Transition Deployment Addressing incompatibilities of cloud solution features during multi-cloud deployment Semantic Interoperability in Multi-Clouds Adoption of the FCLOUDS framework for semantic interoperability in multi-clouds
@@
S32 Fog Computing Deployment Addressing fog computing hardware and container deployment time challenges in multi-cloud environment Reorganized Container Images and Docker Deployment Reorganization of container images and modification of the Docker deployment process
S33 Fog Device Deployment Managing fog device availability and services during multi-cloud container deployment On-Demand Fog and Microservices Deployment Enabling on-demand creation of fogs and deployment of microservices using Docker and Kubeadm
S36 Fog Computing Container Deployment Overcoming low rates of container deployment, especially in fog computing environments FogDocker for Container Deployment Deployment of containers using FogDocker after essential file downloads
S39 Distributed Compute Node Deployment Effectively managing containerized applications across distributed compute nodes in multi-cloud environment Label-Based Scheduling Strategy Development of a label-based scheduling strategy
S40 Kernel Security Deployment Addressing security risks related to kernel sharing, isolation, and OS-level security during multi-cloud container deployment Enhanced Container Security Enhancing container security from the host OS level
@@
S27 API Management Deployment Effectively managing different cloud providers’ APIs/interfaces for multi-cloud container deployment DevOps Approach for Multi-Cloud Applications Adoption of a DevOps approach for managing the lifecycle of multi-cloud applications
S29 Web Service Deployment Ensuring correct and efficient deployment of web services across various computing environments Kubernetes-Based Containerized Deployment Deployment of containerized environments using Kubernetes
S51 Complex Application Deployment Overcoming complexities associated with application deployment in multi-cloud environment Aneka Platform for Distributed Applications Adoption of the Aneka platform for programming and deploying distributed applications
@@
S53 High-availability Deployment Ensuring high availability and preventing outages during deployment in multi-cloud environment Middleware Support for High Availability in Multi-Cloud PaaS Introduction of novel Middleware Support for High Availability in multi-cloud PaaS
S69 Complex Workflow Deployment Utilizing multiple cloud providers for deploying complex application workflows CloudSME Simulation Platform Utilization of the CloudSME Simulation Platform for access to various resources and workflow support
@@
S84 Business Process Provisioning Deployment Orchestrating multi-cloud business process provisioning during deployment Multi-Cloud Service Orchestration Frameworks Adoption of multi-cloud service orchestration frameworks
@@ 3.8. Monitoring Challenges and Solution Framework (RQ4 and RQ5)
Table 14 provides an extensive overview of the monitoring challenges and solution framework for containerized applications in multi-cloud environment. It encompasses diverse challenges related to monitoring containerized applications in the complex multi-cloud landscape. Each category in the figure targets specific monitoring aspects and offers solutions to address these challenges.
Performance, Consistency, and Variability Monitoring: This category emphasizes sustaining performance monitoring in dynamically varying cloud environments for containerized applications. It tackles the challenge of ensuring consistent performance monitoring and proposes solutions like performance variation monitoring and benchmarking for optimized performance.
Resource and Infrastructure Monitoring: In this category, the focus is on monitoring resource and infrastructure management for containerized applications in multi-cloud environment. Challenges include network communication, system image and configuration monitoring, and dynamic resource management. Solutions involve service mesh implementation, Docker server deployment, and using Platform-as-a-Service (e.g., Aneka PaaS) for Cloud Computing.
Security, Compliance, and Trust Monitoring: This category addresses the vital aspects of monitoring security, compliance, and trust management for containerized applications in multi-cloud settings. Challenges include accessing provider host information, continuous security control, and multi-cloud security compliance. Solutions include ML-based monitoring and security assurance platforms in the form of a SaaS (e.g.,MUSA), which support multi-cloud application runtime security control and transparency to enhance user trust.
Application Optimization and Adaptability Monitoring: The challenges of this category revolve around monitoring application optimization and adaptability in multi-cloud environment. Challenges include real-time application monitoring, microservices optimization, application behavior, download execution, and design and state monitoring. To address these challenges, solutions encompass the Dynamic Resource Allocation (DRA) framework, Microservices-based Cloud Application-level Dynamic Orchestrator (MiCADO) optimization, Topology and Orchestration Specification for Cloud Applications (TOSCA) event modeling, FogDocker implementation, and adopting a cloud modeling framework.
System Complexity and Standardization Monitoring: This category focuses on addressing system complexity and promoting standardization in monitoring practices across multi-cloud environment. Challenges include proprietary monitoring solutions and resource isolation mechanisms in the Linux kernel. Solutions encompass pervasive monitoring and identifying in-container leakage channels.
Multi-cloud Coordination and Integration Monitoring: In this category, the challenges revolve around coordinating and integrating monitoring data across multiple cloud environments. Challenges include multi-cloud monitoring challenges, cross-level monitoring improvement, and integrated multi-cloud management monitoring. Solutions encompass self-healing techniques, business process monitoring, and integration with tools like Gnocchi.

### v2

@@ 3.7. Deployment Challenges and Solution Framework (RQ4 and RQ5)
(Sabbioni et al., 2020) Multi-cloud PaaS Deployment Lightweight Proxies for PaaS Integration
(Bou Ghantous and Gill, 2021) IoT Application Deployment DRA Framework and CI Broker for Multi-Cloud Deployment
(Ramesh et al., 2023) Data Transfer and Service Compatibility FaaS and Storage for Multi-Cloud Deployment
(Nguyen and Chien, 2023) Heterogeneous Cloud and Edge Integration RBAM for Cloud-Edge Integration
(Tiwari and Sharma, 2023) Private Cloud Integration Problem CLI Tool for Serverless Setup
@@
(Slominski et al., 2015) Legacy Code Migration Deployment Reusable Architectural Pattern with Docker and Cloudant
(Wang et al., 2020) Tenant Performance Deployment Multi-Tenant and Multi-Instance Hybrid Deployment
(Trihinas et al., 2015) Monitoring Deployment Platform-Independent Monitoring System
(Struhár et al., 2024) Resource Contention and Performance Problem Linux Patch for Dynamic CPU Adaptation
(Horchulhack et al., 2024) QoS Degradation Detection Problem Kubernetes with Performance Monitoring Tools
(Marques et al., 2024) Monitoring Integration in Hybrid Cloud Issue Custom Scaling with Kubernetes
(Kataru et al., 2023) AWS Deployment Performance Consistency AWS Model Comparison for Performance
(Benmerar et al., 2023) Cloud-Edge Deployment Performance Issue AI-driven Orchestration for Dynamic Resource Management
@@
(Barna et al., 2017) Autonomic System Deployment Self-Tuning Performance Model and Autonomic Management
(Fadda et al., 2019) Microservice Deployment Metrics and Requirements Framework for Microservices
(Pop et al., 2016) Multi-cloud Application Configuration Deployment Runtime Environment with Object Store and Artifact Repository
(Quint and Kratzke, 2018) Elastic Container Platform Dependency Deployment Separation of Elastic Platform and Cloud Application
(Hossain et al., 2023) Microservices Deployment in Edge Environments Docker, CI/CD, Service Mesh for Microservices
@@
(Carpio et al., 2020) Inter-edge Bandwidth Deployment Modular Edge Cloud Computing Architecture
(Shi et al., 2020) Network Latency Deployment VM Selection for Composite Applications
(Dreibholz et al., 2019) Latency Reduction Deployment Edge-Cellular Hybrid Infrastructure Provisioning
(Pellegrini et al., 2017) Vendor Lock-in Prevention Deployment Interoperability Layer Above Cloud Infrastructure
(Lee and Nam, 2023) Dynamic Network Policy Deployment Problem KUNERVA for Network Policy Automation
@@
(Pham et al., 2015) ACS Compatibility Deployment Scalable Multi-Cloud Deployment Framework
(Sheridan et al., 2017) Manual Intervention Risk Deployment Automated Security Measures at Deployment
(Wurster et al., 2018) Multi-provider Deployment TOSCA-Based Deployment Modeling Approach
(Challita et al., 2017) Cloud Provider Transition Deployment Semantic Interoperability in Multi-Clouds
(da Silva et al., 2023) Automation Tools Deployment Issue Accurate Cost and Resource Tracking
(Santos et al., 2023) Regional Infrastructure Deployment Consistency Issue AWS Wavelength for Edge Cloud Integration
@@
(Al-shammari and Alwan, 2018) Fog Computing Deployment Reorganized Container Images and Docker Deployment
(Sami and Mourad, 2020) Fog Device Deployment On-Demand Fog and Microservices Deployment
(Lorenzo et al., 2019) Fog Computing Container Deployment FogDocker for Container Deployment
(Altran et al., 2022) Distributed Compute Node Deployment Label-Based Scheduling Strategy
(Bélair et al., 2019) Kernel Security Deployment Enhanced Container Security
(Barletta et al., 2024) Criticality Constraints in Kubernetes Deployment Kubernetes for Critical Deployments
(Centofanti et al., 2024) Power Tool Compatibility Problem Docker, Kubernetes for Power Monitoring Integration
(Ramesh et al., 2023) Data Transfer and Service Compatibility FaaS and Storage for Multi-Cloud Deployment
@@
(Alonso et al., 2019) API Management Deployment DevOps Approach for Multi-Cloud Applications
(Gupta et al., 2021) Web Service Deployment Kubernetes-Based Containerized Deployment
(Buyya and Barreto, 2015) Complex Application Deployment Aneka Platform for Distributed Applications
@@
(Acquaviva et al., 2017) High-availability Deployment Middleware Support for High Availability in Multi-Cloud PaaS
(Kiss, 2018) Complex Workflow Deployment CloudSME Simulation Platform
@@
(Kritikos et al., 2019) Business Process Provisioning Deployment Multi-Cloud Service Orchestration Frameworks
(Roy and Kalafatis, 2023) Robotic App Deployment Challenge Hybrid Edge-Cloud Architecture
(Sofia et al., 2023) Dynamic Adaptation Deployment Issue Context-aware Resource Allocation Strategies
Multi-cloud Deployment Coordination and Integration This category identifies and classifies the challenges and solutions related to coordination and integration during the multi-cloud deployment phase. Of the 121 reviewed studies, 5 addressed these challenges. One such challenge is Multi-cloud PaaS Deployment, which seeks to overcome high entry barriers for deploying a PaaS infrastructure across multiple clouds. The proposed solution is Lightweight Proxies for PaaS Integration, which allows for the seamless integration of different PaaS services through lightweight proxies. One other challenge includes IoT Application Deployment, which presents issues when deploying IoT applications on multi-cloud platforms. The solution for this challenge is the DRA Framework and CI Broker for Multi-Cloud Deployment, which ensure smooth deployment and resource coordination across multiple clouds. Additionally, Data Transfer and Service Compatibility issues are managed using FaaS and storage services, ensuring that data transfer and compatibility between different cloud providers are not problematic. RBAM supports heterogeneous cloud and edge integration, offering frictionless management and operational flexibility for cloud-edge configurations. Lastly, private cloud integration is facilitated by the CLI tools that enable easy integration and the establishment of serverless environments.
Monitoring and Scalability Challenges category identifies and classifies various issues and solutions related to monitoring, performance, and scalability in the context of multi-cloud container deployment. Of the 121 reviewed studies, 8 specifically addressed these issues. One challenge is Legacy Code Migration Deployment, which involves managing the migration of legacy code, tenant engine separation, and composition modeling during container deployment across multiple clouds. The proposed solution is a Reusable Architectural Pattern with Docker and Cloudant, which enables the reuse of an architectural pattern using Docker for containerization and Cloudant for the persistence layer. Another challenge is Tenant Performance Deployment, which focuses on managing tenant service performance and competition during multi-cloud containerized deployments. A proposed solution is the Multi-tenant and Multi-instance Hybrid Deployment scheme, based on container technology, which improves performance efficiency across tenants. A critical challenge is the monitoring of deployments, where existing tools have limitations in multi-cloud environment. The solution involves implementing JCatascopia, a platform-independent and interoperable monitoring system, which enhances the monitoring of multi-cloud setups and helps mitigate these challenges. Additional challenges include Resource Contention and QoS Degradation. The solutions include Dynamic CPU Adaptation through Linux patches, along with Kubernetes and performance monitoring tools. These measures ensure optimal resource usage and maintain performance consistency across cloud environments.
Microservices Architecture and Containerization category classifies the challenges and solutions related to microservices architecture and containerization in multi-cloud environment. Of the 121 reviewed studies, 5 explicitly addressed challenges in this area. One such challenge involves the deployment of multilayer and multitier Web architectures in multi-cloud environment, referred to as Autonomic System Deployment. The proposed solution is a Self-Tuning Performance Model and Autonomic Management, which introduces a self-tuning performance model along with an autonomic management system to optimize deployment and management. Another challenge focuses on Microservice Deployment, which explores how microservice architectures can be deployed in multi-cloud environment. The proposed solution is a Metrics and Requirements Framework for Microservices, emphasizing the need for a structured framework to develop requirements and relevant metrics for deploying microservice-based applications. For Multi-cloud Application Configuration, the challenges involve deployment, configuration, and operation across multi-cloud environment. The one solution involves deploying a Runtime Environment with an Object Store and Artifact Repository to effectively manage configurations across different clouds. The Elastic Container Platform Dependency Deployment solution addresses the challenge of vendor lock-in and dependency on specific container platforms. The solution is the Separation of Elastic Platform and Cloud Application Definitions, which separates platform definitions from cloud application definitions to increase flexibility of application deployment. Additionally, deploying microservices in an edge environment can help resolve integration and communication issues in heterogeneous edge computing environments. The use of Docker, CI/CD Pipelines, and Service Mesh for Microservices enables the automated and secure deployment and communication of microservices.
Network Connectivity and Hybrid Cloud Integration category classifies the challenges and solutions related to ensuring proper network connectivity in multi-cloud deployments and hybrid cloud integration. Out of the 121 studies reviewed, 5 specifically address these issues. One challenge, Inter-edge Bandwidth Deployment, deals with interconnecting distributed localities and managing bandwidth at the edge. The solution is a Modular Edge Cloud Computing Architecture, using containerization for better bandwidth management. A key challenge is Network Latency Deployment, which focuses on minimizing latency during multi-cloud deployments. The proposed solution is VM Selection for Composite Applications, optimizing virtual machine selection to reduce latency. Latency Reduction Deployment involves improving end-to-end performance through an Edge-Cellular Hybrid Infrastructure, combining edge and cellular networks. To address Vendor Lock-in, an Interoperability Layer Above Cloud Infrastructure enables smoother transitions between cloud providers. Finally, Dynamic Network Policy Deployment tackles the challenge of adapting network policies to container workloads, with the solution being Automation of Network Policies using tools like KUNERVA.
Cloud Deployment Constraints and Challenges category identifies constraints and challenges in multi-cloud deployment, with 6 out of 121 studies addressing these issues. One challenge, ACS Compatibility Deployment, involves ensuring container compatibility with existing ACS systems, and the solution is a Scalable Multi-Cloud Deployment Framework. Other challenges include Manual Intervention Risk Deployment, which addresses risks from manual intervention. The solution is Automated Security Measures at Deployment to reduce vulnerabilities. The challenge of Multi-Provider Deployment involves managing multiple cloud providers. The proposed solution is the TOSCA-Based Deployment Modeling Approach, which provides a unified framework. The challenge of Cloud Provider Transition Deployment lies in addressing feature incompatibilities. The proposed solution is Semantic Interoperability in Multi-Clouds, which ensures smooth transitions between providers. Automation Tools Deployment challenge is solved with tools like Jenkins and Terraform for accurate cost and resource management.
Infrastructure Provisioning and Container Deployment Challenges category classifies the challenges and solutions associated with infrastructure provisioning and container deployment in multi-cloud settings. Of the 121 reviewed studies, 8 focused on these challenges. One such challenge is Fog Computing Deployment, which deals with issues related to fog computing hardware and container deployment time. The proposed solution is Reorganized Container Images and Docker Deployment, suggesting that container images should be reorganized, and the Docker deployment process is modified to handle fog computing more effectively, enabling better resource management and optimization. Another challenge is Fog Device Deployment, which focuses on managing fog device availability during multi-cloud container deployments. The proposed solution is On-Demand Fog and Microservices Deployment, allowing for the on-demand creation of fog nodes and the on-demand deployment of microservices via Docker and Kubeadm. Distributed Compute Node Deployment involves challenges related to containerized application management in distributed compute nodes. Kernel Security Deployment addresses the security risks of kernel sharing and isolation. The proposed solution includes a Label-Based Scheduling Strategy and Enhanced Container Security for secure and efficient deployment.
Application Deployment Workflow and Orchestration Challenges category classifies the challenges and solutions related to application deployment workflows and orchestration in multi-cloud environment. Of the 121 reviewed studies, 7 focused on these challenges. One such challenge, API Management Deployment, managing the APIs and interfaces provided by different cloud providers in multi-cloud container deployments. The proposed solution is a DevOps Approach for Multi-Cloud Applications, which suggests using DevOps to manage application lifecycles in multi-cloud environment. Another challenge, Web Service Deployment, involves ensuring proper and efficient deployment in heterogeneous computing environments. The solution is Kubernetes-Based Containerized Deployment, which uses Kubernetes to efficiently manage and deploy containerized environments. Complex Application Deployment and High-availability Deployment focus on deploying complex applications while ensuring high availability. The solutions include the Aneka Platform for Distributed Applications and Multi-cloud PaaS Middleware Support for High Availability, which ensure the availability and distribution of applications across multi-cloud environment.
@@ 3.8. Monitoring Challenges and Solution Framework (RQ4 and RQ5)
Table LABEL:tab:monitoring-solutions provides an extensive overview of the monitoring challenges and solution framework for containerized applications in multi-cloud environment. It encompasses diverse challenges related to monitoring containerized applications in the complex multi-cloud landscape. Each category in the figure targets specific monitoring aspects and offers solutions to address these challenges.
Performance, Consistency, and Variability Monitoring category identifies the challenges and solutions that ensure consistent performance monitoring in dynamically changing multi-cloud environment. Nine of the reviewed studies addressed these challenges. Performance Monitoring focuses on maintaining consistent performance tracking of applications. The solution, Performance Variation Monitoring, uses benchmarking systems to track and analyze runtime changes in performance, ensuring optimized performance throughout. Another challenge is Benchmarking for Performance and Cost Efficiency, which balances performance against cost. The solution, SBDO Optimization, optimizes resource utilization in real time to maximize efficiency and minimize costs. A related challenge, Website Performance Monitoring, is managed through Kubernetes Monitoring, which ensures uptime and performance across widely distributed cloud environments. Monitoring User Experience is handled by collecting CSP QoS Metrics, allowing cloud service providers to gather Quality of Service data, anticipate variations in performance, and manage the overall user experience effectively. Lastly, low-latency performance optimization in multi-cloud systems is addressed by Tail Latency Prediction, which uses predictive models to minimize network delays and ensure smooth application performance.
Resource and Infrastructure Monitoring catagory classifies the challenges and corresponding solutions related to monitoring resource and infrastructure management for containerized applications in multi-cloud environment. Of the reviewed studies, 8 specifically addressed these challenges. One key challenge is Network Communication Monitoring, which involves managing service-to-service communication across cloud networks. The proposed solution is the implementation of a Service Mesh, which enhances visibility and security for network services. Another challenge is System Image and Configuration Monitoring, which tracks software updates, configuration changes, and performance across platforms. The solution proposed is Docker Server Deployment, which allows for easy updates to the containerized system. For Vcluster and Framework Management Monitoring, the solution is the implementation of LXC (Linux Containers), which manages application isolation and performance using lightweight containers. Multi-Cloud Resource Utilization Monitoring involves tracking the performance of resources, such as compute, storage, and network, used across multiple clouds. The solution uses Prometheus and Grafana for monitoring resource usage and availability, with an emphasis on optimization. Similarly, Dynamic Resource Management Monitoring solutions, such as the Aneka Platform, dynamically add or release resources based on demand. The challenge of cost tracking in resource utilization across multiple cloud environments, such as AWS and Azure, is addressed by the COSTA system, which monitors both resource utilization and cost efficiency. Finally, Prometheus is used for Real-Time Dynamic Resource Monitoring, allowing real-time data collection and analysis to optimize performance based on resource usage fluctuations.
Table 14. Identified monitoring challenge and solution for containerized applications in multi-cloud environment
ID Challenge Solution
Category 1: Performance Monitoring Consistency and Variability
(Almeida et al., 2014) Performance Monitoring Performance Variation Monitoring
(Jha et al., 2019) Benchmarking Performance Variation and Cost Efficiency Monitoring SDBO Optimization
(Gupta et al., 2021) Website Performance and Availability Monitoring Kubernetes Monitoring
(Pandey et al., 2022) User Experience Variability Monitoring CSP QoS Metrics Collection
(Dreibholz et al., 2019) Independent Performance Metrics Monitoring Performance Evaluations
(Bhamare et al., 2015) Comprehensive Network Performance Measurement Monitoring Server Selection Scheme
(Rahman and Lama, 2019) Microservice Performance Degradation Monitoring Tail Latency Prediction
(Struhár et al., 2024) Container Performance Metrics Tracking CLM and OSLM for Resource Adjustments
(Horchulhack et al., 2024) Container Performance and Resource Usage Monitoring Metric and Container Monitoring Tools
(Hossain et al., 2023) Microservice Performance Monitoring Challenge Prometheus, Grafana, Kubernetes for Monitoring
(Ramesh et al., 2023) Cross-Cloud Service Performance Monitoring FaaS and Storage Services for AI Monitoring
(Nguyen and Chien, 2023) Stream Processing Performance Monitoring RBAM for Real-Time Performance Insights
(Kataru et al., 2023) AWS Application Performance Monitoring AWS CloudWatch for Performance Monitoring
(Alonso et al., 2023a) Multi-Cloud Performance Visibility Issue Monitoring Tools for Multi-Cloud Visibility
(Harzenetter et al., 2023) Cross-Technology Performance Monitoring Issue TOSCA-based Workflows for Monitoring Applications
(Huang and Wang, 2023) Industrial Robot System Monitoring Challenge Real-Time System Performance Tracking Tools
Category 2: Resource and Infrastructure Management Monitoring
(Gattobigio et al., 2022) Network Communication Monitoring Service Mesh Implementation
(Verdugo et al., 2017) System Image and Configuration Monitoring Docker Server Deployment
(Cui et al., 2016) Vcluster and Framework Management Monitoring LXC Implementation
(Pahl, 2015) Multi-Cloud Resource Utilization Monitoring Prometheus and Grafana Monitoring
(Buyya and Barreto, 2015) Dynamic Resource Management Monitoring Aneka Platform
(da Silva et al., 2023) Cost and Resource Usage Monitoring COSTA for Cost and Resource Monitoring
(Dogani et al., 2023) Real-Time Dynamic Resource Monitoring Prometheus for Real-Time Monitoring
(Wong et al., 2023) Resource Usage and Performance Issue Prometheus for Real-Time Data Collection
(Marques et al., 2024) Real-Time Container Monitoring Prometheus with Forecasting for Resource Management
(Santos et al., 2023) Distributed Cloud Resource Monitoring Issue AWS CloudWatch and Azure Monitor for Insights
(Nagarajan et al., 2023) Resource Allocation and Energy Monitoring Issue Tools for Resource, Energy, and SLA Monitoring
Category 3: Security, Compliance, and Trust Management Monitoring
(Ramos et al., 2021) Provider Host Information Access Monitoring ML Monitoring
(Rios et al., 2017) Continuous Security Control Monitoring MUSA Platform
(Casola et al., 2017) Multi-Cloud Security Compliance Monitoring MUSA Security Compliance
(Acquaviva et al., 2017) Service Continuity and Data Privacy Monitoring NoMISHAP Implementation
(Wahab et al., 2016) Trust Relationship and Network Contact Monitoring Trust Bootstrapping
(Fugkeaw, 2023) Access and Authentication Monitoring Issue Blockchain for Secure Access Event Logging
(Wang et al., 2023) Malware Behavior and System Monitoring DockerWatch for Container Behavior Monitoring
(Neeraj et al., 2023) Log Data Integrity Monitoring Issue Blockchain-based Logs for SLA Monitoring
Category 4: Application Optimization and Adaptability Monitoring
(Ferry et al., 2014) Adaptive System Design and State Monitoring Cloud Modelling Framework
(Lorenzo et al., 2019) Partial Download Execution Monitoring FogDocker Implementation
(Wurster et al., 2018) Event-Driven Application Behavior Monitoring TOSCA Event Modeling
(Guerrero et al., 2018a) Microservice Optimization Monitoring MiCADO Optimization
(Bou Ghantous and Gill, 2021) Real-Time Application Monitoring DRA Framework
(Sofia et al., 2023) Container Application Adaptability Monitoring Advanced Monitoring with Orchestration Integration
(Petrovski and Gusev, 2024) Serverless and Containerized Monitoring Challenge AWS Tools for Performance Monitoring
Category 5: System Complexity and Standardization Monitoring
(Sabbioni et al., 2020) Proprietary Monitoring Solution Standardization Pervasive Monitoring
(Gao et al., 2017) Kernel Resource Isolation Mechanism Monitoring Leakage Defense
(Alonso et al., 2019) Multi-Cloud Application Monitoring Complexity DECIDE DevOps Expansion
(Barna et al., 2017) Monitoring Agent Coordination Autonomic Management
(Sami and Mourad, 2020) Heterogeneous Fog Resource Management Monitoring Kubernetes Utility Architecture
(Baresi et al., 2016) Application Topology and Dependency Monitoring MicroCloud TOSCA Library
(Trihinas et al., 2015) Elastic Cloud Service Monitoring Research JCatascopia Monitoring
(Fadda et al., 2019) Microservice Monitoring Diversity QoM Definition
(Harwalkar et al., 2019b) Compute and Network Resource Utilization Monitoring Time Series Resource Utilization
(Ahmad and Aujla, 2023) Real-Time Data Operations Monitoring Smart Contracts for GDPR Monitoring
(Centofanti et al., 2024) Power Consumption Monitoring Issue Prometheus with Power Monitoring Integration
(Tiwari and Sharma, 2023) Serverless Function Monitoring Complexity CLI Tool for Function Monitoring
Category 6: Multi-cloud Coordination and Integration Monitoring
(Sitaram et al., 2018) Integrated Multi-Cloud Management Monitoring Gnocchi Integration
(Mfula and Nurminen, 2018) Multi-Cloud Monitoring Challenges Self-Healing Techniques
(Kritikos et al., 2019) Cross-Level Monitoring Improvement Business Process Monitoring
(Lee and Nam, 2023) Real-Time Network Policy Monitoring Challenge KUNERVA for Real-Time Network Policy Monitoring
(Sofia et al., 2024) Decentralized System Observability Issue CODECO for Orchestration and Data Observability
Security, Compliance, and Trust Monitoring: category addresses the challenges and solutions related to monitoring security, compliance, and trust management for containerized applications in multi-cloud environment. Out of the 121 reviewed studies, 9 specifically deal with these concerns. One challenge is Provider Host Information Access Monitoring, where limitations in accessing host data are overcome using ML Monitoring, which employs machine learning to track system performance and usage patterns. Another challenge, Continuous Security Control Monitoring, is addressed by the MUSA platform, which enforces and continuously monitors security in real time for containerized applications. Multi-Cloud Security Compliance Monitoring ensures continuous security compliance across multi-cloud environment. This is achieved through the MUSA Security Compliance Solution, a platform for continuous compliance management. The challenges of maintaining data privacy, fault tolerance, and service continuity in multi-cloud environments are addressed under Service Continuity and Data Privacy Monitoring. The NoMISHAP Implementation is proposed to handle these monitoring issues. Additionally, Trust Bootstrapping establishes trust through a chain of endorsement and decision tree classification. Access and Authentication Monitoring uses blockchain technology to enable secure and transparent logging of authentication activities. For Malware Behavior and System Monitoring, DockerWatch monitors container behavior with minimal performance overhead.
Application Optimization and Adaptability Monitoring: This category identifies and discusses challenges related to optimizing and adapting applications in multi-cloud environment. Out of the 121 reviewed studies, 8 specifically addressed these challenges with corresponding solutions. Key challenges include Real-Time Application Monitoring, which focuses on taking immediate corrective actions using the DRA Framework, and Microservice Optimization Monitoring, which enhances monitoring across multiple clouds through MiCADO Optimization. Other challenges include Partial Download Execution Monitoring, which ensures the correct execution of partially downloaded files. The proposed solution for runtime accuracy is the FogDocker Implementation. Another challenge, Event-Driven Application Behavior Monitoring, involves monitoring event-driven application behavior. The TOSCA Event Modeling solution efficiently handles application events. For Adaptive System Design and State Monitoring, continuous design and dynamic system adaptability are managed using the Cloud Modelling Framework. Additionally, the challenge of Container Application Adaptability Monitoring is addressed through advanced orchestration techniques, which provide real-time views of distributed applications. The solution involves using AWS tools to ensure high performance for different deployment models, addressing challenges in Serverless and Containerized Monitoring. AWS tools are recommended to ensure high performance for different deployment models, addressing challenges in Serverless and Containerized Monitoring.
System Complexity and Standardization Monitoring cateogy identifies and addresses the challenges of managing system complexity and standardization in monitoring solutions across multi-cloud environment. Of the 121 studies reviewed, 4 specifically focused on these challenges and proposed corresponding solutions. A key challenge is the standardization of proprietary monitoring solutions, for which the proposed solution is Pervasive Monitoring. Pervasive Monitoring focuses on implementing pervasive monitoring techniques for containerized applications within multi-cloud PaaS platforms. Another significant challenge is Kernel Resource Isolation Mechanism Monitoring, with the proposed solution being Leakage Defense. This solution addresses in-container leakage channels by implementing a two-stage isolation mechanism in Linux-based systems. Additionally, the challenge of Multi-Cloud Application Monitoring Complexity is addressed by the DECIDE DevOps Expansion, which extends the framework for analyzing multi-cloud containerized applications. The Autonomic Management Framework further enhances monitoring through Monitoring Agent Coordination, improving the coordination of monitoring agents using a performance model and adaptive actions for better system behavior within containerized applications.
Multi-cloud Coordination and Integration Monitoring category identifies and addresses the challenges related to coordinating and integrating monitoring data across multiple cloud environments. Out of the 121 studies reviewed, 5 focused on these challenges. One key challenge is Multi-cloud Monitoring Improvement, and the proposed solution is the use of Self-Healing Techniques to autonomously detect and correct issues within cloud infrastructures. Another challenge is Cross-Level Monitoring, which is addressed through Business Process Monitoring Integration to ensure that business processes are effectively monitored across multiple cloud platforms. Lastly, Integrated Multi-cloud Management Monitoring requires advanced solutions, such as Integration with Gnocchi, to efficiently gather and monitor metrics from diverse cloud resources.